140830 LML Observations

Isaac arrived at LML at 10:20 am. The doors to the collector had been deployed and a sample had been collected in the sample jar. The sample measured approximately 50 ml. The sample was relatively clear with a small amount of sediment at the bottom. The sample was then removed from the collector and placed in a small labeled bag. The collector was then dissembled and cleaned with 409 and DI water. After the collector was reassembled a blank was then taken which measured approximately 200ml. The collector was then reset for the next fog event.

Leave a Reply

Your email address will not be published. Required fields are marked *